# A Weak Penalty Neural ODE for Learning Chaotic Dynamics from Noisy Time Series

### v1

@@ Abstract
Accurate forecasting of complex high-dimensional dynamical systems from observational data is essential for several applications across science and engineering. A key challenge, however, is that real-world measurements are often corrupted by noise, which severely degrades the performance of data-driven models. Particularly, in chaotic dynamical systems, where small errors amplify rapidly, it is challenging to identify a data-driven model from noisy data that achieves short-term accuracy while preserving long-term invariant properties. In this paper, we propose the use of the weak formulation as a complementary approach to the classical strong formulation of data-driven time-series forecasting models. Specifically, we focus on the neural ordinary differential equation (NODE) architecture. Unlike the standard strong formulation, which relies on the discretization of the NODE followed by optimization, the weak formulation constrains the model using a set of integrated residuals over temporal subdomains. While such a formulation yields an effective NODE model, we discover that the performance of a NODE can be further enhanced by employing this weak formulation as a penalty alongside the classical strong formulation-based learning. Through numerical demonstrations, we illustrate that our proposed training strategy, which we coined as the Weak-Penalty NODE (WP-NODE), achieves state-of-the-art forecasting accuracy and exceptional robustness across benchmark chaotic dynamical systems.
@@
However, training NODEs remains challenging in practice. Standard Neural ODE training uses the adjoint sensitivity method [30, 31, 26], which, despite its memory efficiency, incurs high computational cost from repeated ODE forward and backward solves. Moreover, backward integration is often unstable, with exploding or vanishing gradients [32, 33, 34, 35], especially in stiff systems [36, 37]. Additionally, vanilla NODEs do not inherently guarantee stable long-horizon rollouts; achieving robustness typically requires additional architectural constraints or specialized training techniques [33, 32].
@@
In the following sections, WP-NODE demonstrates its versatility and efficacy across several chaotic systems, including the 3-dimensional Lorenz-63 system, the 40-dimensional Lorenz-96 system, and the 64-dimensional Kuramoto–Sivashinsky equation. Our results show that this weak penalty formulation consistently improves both short-term predictive accuracy and long-term statistical consistency, offering a new perspective on robust data-driven modeling of complex dynamical systems.
@@ I Results
Note that with 20% noise, WP-NODE utilizes 5 rollouts, which is found to be more effective for short-time prediction compared to lower-noise cases, where only a single rollout is used. This longer rollout strategy significantly improves VPT on the KS system, potentially because exposure to longer temporal contexts during training helps the model better distinguish signal from noise. In Supplementary Note 6, we provide the details of how many rollouts are used for each case in Table 1.
@@ II Conclusion and Discussion
This work introduced a weak penalty in training NODE, named WP-NODE, for learning chaotic dynamical systems from noisy data. Particularly, the proposed approach is to mitigate the lack of prediction skill of the standard NODE trained with a strong form formulation when the training data are corrupted by noise. Furthermore, the strong form NODE model’s long-term statistical prediction accuracy is sensitive to the choice of ODE integrator. While weak-form supervision alone achieves relatively high predictive horizons, it lacks statistical fidelity. We found that the proposed WP-NODE effectively reconciles both extended prediction time and robustly preserves long-term statistical structure. Across three chaotic systems, L63, L96, and KS, WP-NODE consistently delivers accurate short-term forecasts and preserves faithful long-term statistics, even under substantial observational noise.
@@
Beyond efficiency, WP-NODE benefits from its combined weak–strong structure: the weak formulation naturally mitigates overfitting compared to pure pointwise losses, while the strong-form rollout enhances numerical stability by explicitly solving the underlying ODE. In many cases, the weak form alone trains successfully but fails at inference time due to exponentially exploding dynamics, making the weak penalty approach particularly effective for learning from noisy data.
However, the implemented weak formulation adopts predefined subdomain sizes and fixed test functions, which may influence performance in certain settings. While extensive ablation studies confirm robustness across a range of configurations, further improvements may be possible through adaptive subdomain partitioning, learnable test functions, or task-specific architectural tuning—particularly for more complex or higher-dimensional systems.
@@ III Method
This construction ensures vanishing at the boundaries $(\phi_{p}(\pm 1)=0)$, allows smoothness to be tuned by the integer $p$, and, being polynomial, permits semi-analytical evaluation of the weak-form integrals. Detailed derivations are provided in the Supplementary Note 4.
@@
$\mathcal{L}_{\text{weak}}(\theta)=\big|\mathbf{u}^{\top}\mathbf{w}_{\text{lhs}}+\mathbf{f}^{\top}\mathbf{w}_{\text{rhs}}\big|^{2},$ (10)
where $\mathbf{f}$ denotes the vector of neural network evaluations at the nodal points $\mathbf{u}$ in each subdomain, i.e., $\mathbf{f}=\bigl[f(u_{0},t_{0};\theta),f(u_{1},t_{1};\theta),\dots,f(u_{M-1},t_{M-1};\theta)\bigr]^{\top}$. This formulation naturally generalizes to higher-dimensional or vector-valued data via batched dot products across dimensions. Detailed derivations, including the piecewise linear interpolation, the explicit form of the basis functions $\ell_{i}(s)$, and the construction of the weights, are provided in the Supplementary Note 5.
@@
$\mathcal{L}=\mathcal{L}_{\text{weak}}+\lambda,\mathcal{L}_{\text{strong}},$ (11)
@@
Dataset and Training. Training samples were generated using a sliding-window approach, where each windowed segment defines a short trajectory of fixed length, enabling consistent training across strong, weak, and WP-NODE formulations. Data is scaled using a MinMax scaler fitted on the training data. All neural network models used shallow neural networks (1–2 layers) with GELU activations and were trained with the Adam optimizer. A learning rate scheduler and early stopping based on validation loss are applied to ensure efficient convergence. For inference and reference solution generation, the adaptive dopri5 solver is employed in double precision with a strict numerical tolerance. Full details on all hyperparameters (e.g., subdomain sizes, batch sizes, learning rates, etc.) and other configurations are provided in Supplementary Note 6.
@@
Author contributions J.H. and R.M. supervised the study. All authors contributed to the initial concept and machine learning methodology. X.L. carried out model development, implementation, training, and analysis. All authors contributed to writing the manuscript.
@@ III Supplementary Note 3: Lorenz 96 analysis
IV Supplementary Note 4: Polynomial Test Functions
@@
V Supplementary Note 5: Piecewise Linear Interpolation and Decoupling of Integral
@@
VI Supplementary Note 6: Implementation and training details
@@
Supplementary Table 1: Number of rollouts used during WP-NODE training for each case in Table 1.

### v2

@@ Abstract
Accurate forecasting of complex high-dimensional dynamical systems from observational data is essential for several applications across science and engineering. A key challenge, however, is that real-world measurements are often corrupted by noise, which severely degrades the performance of data-driven models. Particularly, in chaotic dynamical systems, where small errors amplify rapidly, it is challenging to identify a data-driven model from noisy data that achieves short-term accuracy while preserving long-term invariant properties. In this paper, we propose the use of the weak formulation as a complementary approach to the classical strong formulation of data-driven time-series forecasting models. Specifically, we focus on the neural ordinary differential equation (NODE) architecture. Unlike the standard strong formulation, which relies on the discretization of the NODE followed by optimization, the weak formulation constrains the model using a set of integrated residuals over temporal subdomains. While such a formulation yields an effective NODE model, we discover that the performance of a NODE can be further enhanced by employing this weak formulation as a penalty alongside the classical strong formulation-based learning. Through numerical demonstrations, we illustrate that our proposed training strategy, which we coined as the Weak-Penalty NODE (WP-NODE), achieves state-of-the-art forecasting accuracy and exceptional robustness across benchmark chaotic dynamical systems and real-world climate dataset.
@@
However, training NODEs remains challenging in practice. Standard NODE training uses the adjoint sensitivity method [30, 31, 26], which, despite its memory efficiency, incurs high computational cost from repeated ODE forward and backward solves. Moreover, backward integration is often unstable, with exploding or vanishing gradients [32, 33, 34, 35], especially in stiff systems [36, 37]. Additionally, vanilla NODEs do not inherently guarantee stable long-horizon rollouts; achieving robustness typically requires additional architectural constraints or specialized training techniques [33, 32].
@@
In the following sections, WP-NODE demonstrates its versatility and efficacy across four chaotic systems, including the 3-dimensional Lorenz-63 system, the 40-dimensional Lorenz-96 system, the 64-dimensional Kuramoto–Sivashinsky system, and a real-world climate dataset, ERA5. Our results show that this weak penalty formulation consistently improves both short-term predictive accuracy and long-term statistical consistency, offering a new perspective on robust data-driven modeling of complex dynamical systems.
@@ I Results
Note that with 20% noise, WP-NODE utilizes 5 rollouts, which is found to be more effective for short-time prediction compared to lower-noise cases, where only a single rollout is used. This longer rollout strategy significantly improves VPT on the KS system, potentially because exposure to longer temporal contexts during training helps the model better distinguish signal from noise. In Supplementary Note 7, we provide the details of how many rollouts are used for each case in Table 1.
@@
Figure 5: NODE predictions performance on real-world ERA5 dataset. a. Short-term (14-day) temperature prediction errors. WP-NODE achieves the lowest error. b. Long-term (10-year) temperature prediction errors. WP-NODE remains stable. c. Climatological statistical prediction for 10 years of WP-NODE across the four ERA5 prognostic variables.
ERA5 atmospheric reanalysis dataset. To evaluate WP-NODE in a high-dimensional, noisy, and strongly coupled real-world setting, experiments are conducted on the ERA5 atmospheric reanalysis dataset with a temporal resolution of 6 hours from January 1, 2000 to December 15, 2009 [60]. The data are regridded to a T30 Gaussian grid and vertically interpolated onto $\sigma$-levels [61], yielding input tensors of four prognostic variables, including near-surface temperature ($t$) and specific humidity ($q$) at $\sigma_{0.95}$, and mid-atmospheric zonal ($u$) and meridional ($v$) winds at $\sigma_{0.51}$. The horizontal resolution of each variable consists of 48 latitude times 96 longitudinal data points. Total Incoming Solar Radiation (TISR) is added as a temporal embedding to capture external radiative forcing [62, 32].
A CNN-based NODE architecture is employed for this spatio-temporal dataset. Following model training, we assess and compare both short- and long-term predictive capacity of WP-NODE against the baseline weak and strong NODEs. In the short-term (Fig. 5a, 14 days), both weak and strong NODEs quickly become unstable and show rapidly increasing errors, whereas WP-NODE consistently achieves the lowest RMSE. Even over a 10-year horizon (Fig. 5b), WP-NODE remains stable, while both baseline NODEs diverge. Figure 5c further evaluates the predicted climatological statistics. For each variable, the temporal average over the 10-year testing period is computed. The zonal mean at each latitude is then obtained by averaging the time-averaged field along the longitude direction. Notice that WP-NODE produces accurate predictions of both the zonal mean temperature and specific humidity. While the prediction for the average zonal and meridional winds is less accurate, WP-NODE is able to capture their trend relatively well. Detailed configurations and additional performance comparisons, including extended 1-year predictions for other prognostic variables, are provided in Supplementary Note 4.
Overall, the ERA5 experiment highlights the robustness of our hybrid training strategy and demonstrates that, even with a simple CNN-based architecture, WP-NODE delivers superior stability, accuracy, and long-term predictive fidelity on real-world climate data, where both weak and strong NODEs deteriorate.
@@ II Conclusion and Discussion
This work introduces a weak penalty in training NODE, named WP-NODE, for learning chaotic dynamical systems from noisy data. Particularly, the proposed approach mitigates the lack of prediction skill of a standard NODE trained with a strong form formulation when the training data are corrupted by noise. Furthermore, the strong form NODE model’s long-term statistical prediction accuracy is sensitive to the choice of the ODE integrator. WP-NODE, by contrast, exhibits greater independence from the specific numerical solver employed, enhancing its robustness across different integration schemes.
While weak-form supervision alone achieves relatively high predictive horizons, it lacks statistical fidelity. We found that the proposed WP-NODE effectively reconciles both extended prediction time and robustly preserves long-term statistical structure. Across four chaotic systems, L63, L96, KS, and a climate forecasting benchmark, WP-NODE consistently delivers accurate short-term forecasts and preserves faithful long-term statistics, even under substantial observational noise. Furthermore, application to the ERA5 climate reanalysis dataset demonstrates that WP-NODE maintains stable and accurate long-term forecasts even with a simple CNN-based architecture, where strong and weak NODE approaches quickly diverge.
@@
Beyond efficiency, WP-NODE benefits from its combined weak–strong structure: the weak formulation naturally mitigates overfitting compared to pure pointwise losses, while the strong-form rollout enhances numerical stability by explicitly solving the underlying ODE. In many cases, the weak form alone trains successfully but fails at inference time due to exponentially exploding dynamics, making our approach particularly effective for learning from noisy data.
We note that the implemented weak formulation adopts predefined subdomain sizes and fixed test functions, which may influence performance in certain settings. While extensive ablation studies confirm robustness across a range of configurations, further improvements may be possible through adaptive subdomain partitioning, learnable test functions, or task-specific architectural tuning, particularly for more complex or higher-dimensional systems.
@@ III Method
This construction ensures vanishing at the boundaries $(\phi_{p}(\pm 1)=0)$, allows smoothness to be tuned by the integer $p$, and, being polynomial, permits semi-analytical evaluation of the weak-form integrals. Detailed derivations are provided in the Supplementary Note 5.
@@
$\mathcal{L}_{\text{weak}}(\theta)=\left\|\mathbf{u}^{\top}\mathbf{w}_{\text{lhs}}+\mathbf{f}^{\top}\mathbf{w}_{\text{rhs}}\right\|^{2},$ (10)
where $\mathbf{f}$ denotes the vector of neural network evaluations at the nodal points $\mathbf{u}$ in each subdomain, i.e., $\mathbf{f}=\bigl[f(u_{0},t_{0};\theta),f(u_{1},t_{1};\theta),\dots,f(u_{M-1},t_{M-1};\theta)\bigr]^{\top}$. This formulation naturally generalizes to higher-dimensional or vector-valued data via batched dot products across dimensions. Detailed derivations, including the piecewise linear interpolation, the explicit form of the basis functions $\ell_{i}(s)$, and the construction of the weights, are provided in the Supplementary Note 6.
@@
$\mathcal{L}=\mathcal{L}_{\text{weak}}+\lambda\mathcal{L}_{\text{strong}},$ (11)
@@
Dataset and Training. Training samples were generated using a sliding-window approach, where each windowed segment defines a short trajectory of fixed length, enabling consistent training across strong, weak, and WP-NODE formulations. Data is scaled using a MinMax scaler fitted on the training data. For the L63, L96, and KS systems, shallow neural networks with one or two layers and GELU activations were employed and trained with the Adam optimizer. A CNN was used to model the multidimensional ERA5 dataset. A learning rate scheduler and early stopping based on validation loss are applied to ensure efficient convergence. For inference and reference solution generation, the adaptive dopri5 solver is employed in double precision with a strict numerical tolerance. For the ERA5 data, a simple Euler step is used. Full details on all hyperparameters (e.g., subdomain sizes, batch sizes, learning rates, etc.) and other configurations are provided in Supplementary Note 4 and 7.
@@
Author contributions J.H. and R.M. supervised the study. X.L., J.H., and R.M. contributed to the initial concept and machine learning methodology. X.L. carried out model development, implementation, training, and analysis. D.C. contributed to the EAR5 dataset processing, training, and visualization. All authors contributed to writing the manuscript.
@@ III Supplementary Note 3: Lorenz 96 analysis
IV Supplementary Note 4: ERA5 analysis
The ERA5 dataset was subjected to min-max normalization, and the data were partitioned chronologically into training and validation sets. The training set contains 11,576 samples that span approximately 7.9 years of data. The validation and test sets contain 4,920 samples that cover the remaining period. Model development and model selection were performed using only the training and validation sets.
The NODE for modeling the ERA5 does not use a simple feed-forward network to parameterize the right-hand side. Instead, the derivative function is implemented by a structured architecture composed of three components in sequence. The encoder compresses the input state and conditioning variables into a latent representation. This latent representation is then processed by multiple residual blocks in the dilated CNN processor, with each block applying a sequence of dilated convolutions to capture multi-scale patterns. Finally, the decoder maps the processed latent representation back to physical space, producing the rate of change for each predicted variable. To correctly model the Earth’s spherical shape, all convolutional operations use cylindrical boundary conditions: circular padding for longitude (to connect East and West) and replicate padding for latitude (to handle the poles).
Since solar forcing varies with time and location, our NODE model accepts conditioning information as an additional input. The temporal integration is performed using the Euler method:
$x_{t+1}=x_{t}+f(x_{t},u_{t},t;\theta)\Delta t,$ (S1)
where $x_{t}$ represents the atmospheric state, $u_{t}$ is the time-dependent input forcing (i.e., solar radiation), $\theta$ denotes the learnable parameters, and $\Delta t$ is the integration time step. The time variable is non-dimensionalized by the data sampling interval (6 hours). Therefore, the integration step size is set to $\Delta t=1$ in the discrete model, representing one step of forward prediction corresponding to a 6-hour physical duration. During training, the strong NODE employs 6 rollout steps, while the weak NODE uses a subdomain size of 8. WP-NODE loss combines both settings above with the penalty $\lambda=1$.
Supplementary Figure 6: Short-term prediction comparison (using RMSE) on the ERA5 dataset over the four prognostic variables. a. Temperature. b. Humidity. c. Zonal component of the horizontal wind velocity. d. Meridional component of the horizontal wind velocity.
Supplementary Figure 7: Long-term prediction comparison (using RMSE) on the ERA5 dataset over the other three prognostic variables (i.e., humidity, zonal, and meridional component of the horizontal wind velocity. a. Comparisons of 14-day predictions. b. 1-year prediction performance of WP-NODE.
In Fig. 6, short-term predictions from WP-NODE are compared with two baseline NODE models for 6-hour and 24-hour forecasts across the four prognostic variables. WP-NODE consistently outperforms both baselines across all variables and time horizons.
In Fig. 7a, the 14-day RMSE evolution is shown for the three prognostic variables not included in the main article (the temperature variable was presented in the main text). The weak and strong baseline NODE models diverge and produce unstable predictions within days, while WP-NODE maintains stable and accurate forecasts throughout the 14-day period. Fig. 7b further demonstrates that WP-NODE can maintain forecast stability over significantly longer time horizons. These results further strengthen the claims in the main article that WP-NODE’s hybrid training strategy delivers superior stability, accuracy, and long-term predictive fidelity in real-world climate forecasting, precisely where conventional NODE approaches fail.
V Supplementary Note 5: Polynomial Test Functions
@@
VI Supplementary Note 6: Piecewise Linear Interpolation and Decoupling of Integral
@@
VII Supplementary Note 7: Implementation and training details
@@
Supplementary Table 1: Number of rollouts used during WP-NODE training for each case in Table 1 of the main article.